# Ye Bai1, Minghan Wang1, Thuy-Trang Vu1

### v1

@@ 3.2 Main Results
Table 2 presents the performance comparison on WikiTQ and TabFact across LLaMA3.3-70B and Qwen2.5-72B. Our proposed method, MAPLE, consistently outperforms all baselines across both datasets and model backbones. On WikiTQ, MAPLE  achieves 74.01% and 73.39% accuracy with respective models, while on TabFact, it reaches 90.66% and 86.02%. These results represent substantial gains of up to +8.26% on WikiTQ and +7.41% on TabFact over the strongest baseline methods.
Compared to recent specialized frameworks like Chain-of-Table and ReAcTable, MAPLE  demonstrates consistent improvements across both datasets. The gains are particularly pronounced on WikiTQ (+5.52% with Qwen2.5-72B), aligning with our framework’s strength in handling compositional reasoning tasks that require progressive refinement. For TabFact, the improvements confirm that our approach remains effective even in binary classification settings. Notably, while baseline methods show varying performance between model backbones, MAPLE  maintains its superiority regardless of the underlying LLM, suggesting that our architecture provides fundamental reasoning advantages independent of specific base model capabilities. For analysis of how table size affects reasoning performance and the impact of multi-round reasoning, see Appendix E.
@@ 3.3 Evaluating Agent Contributions
To assess the contribution of each agent in our framework, we conduct an ablation study using the LLaMA3.3-70B model on WiKiTQ. As shown in Table 3, we incrementally introduce agents into the MAPLE  pipeline. Starting from a baseline that directly predicts answers without reasoning, we observe that adding the Solver agent alone leads to a substantial performance boost of +18.2 points, confirming the effectiveness of our ReAct-style multi-step reasoning. Incorporating the Checker further improves accuracy (+2.1), suggesting that verifying answer quality plays a crucial role in reducing erroneous outputs. Introducing the Reflector yields an additional significant gain (+5.2), highlighting the importance of iterative reflection and error correction. Finally, equipping the system with the Archiver enables long-term memory utilization, resulting in peak accuracy of 74.01. These findings demonstrate the complementary roles of all four agents and validate the cumulative benefit of MAPLE’s modular design.
@@ 3.4 Memory Analysis and System Behavior
As $\delta$ increases from 0.3 to 1.0, total memory size decreases dramatically—from 4,078 to 191 notes for WikiTQ and 1,882 to 78 for TabFact. This demonstrates our selective integration approach in action, preventing memory explosion at low thresholds while becoming increasingly selective at higher thresholds. At $\delta=0.3$, nearly all experiences (93.9%) are added to memory, creating potential redundancy, while at $\delta=1.0$, only highly unique experiences (4.4%) are retained, representing two extremes of the memory gradient phenomenon.
@@
Both datasets exhibit remarkably similar memory dynamics despite their different task characteristics, suggesting that these patterns reflect fundamental properties of knowledge organization rather than dataset-specific artifacts. The consistency across task types provides strong evidence for the robustness of our memory evolution approach.
@@
These findings align with the "approximate learning" theory in cognitive science, which posits that optimal knowledge acquisition occurs when new information is related to existing knowledge that is neither too similar nor too different (Gentner and Smith, 2013). Our empirical results showing peak evolution at moderate distances ($\delta\approx 0.7$) provide computational evidence for this cognitive principle.
@@ Appendix A Related Work
Multi-agent System
Multi-agent systems (MAS) have gained attention for leveraging collective intelligence in complex tasks. Current applications span problem-solving domains such as software development (Li et al., 2023a; Du et al., 2023; Qian et al., 2024; Huang et al., 2024), embodied robotic coordination (Dasgupta et al., 2023; Mandi et al., 2023; Zhang et al., 2024; Yu et al., 2025a), and scientific Debate (Du et al., 2023; Xiong et al., 2023; Tang et al., 2024). World simulation applications include social behavior modeling (Park et al., 2022; Gao et al., 2023; Chen et al., 2025), policy simulation (Xiao et al., 2023; Hua et al., 2024), economic forecasting (Horton, 2023; Li et al., 2023b; Zhao et al., 2024), and gaming (Light et al., 2023; Wang et al., 2023b; Xu et al., 2024). Several MAS frameworks, including MetaGPT (Hong et al., 2024), CAMEL (Li et al., 2023a), and AutoGen (Wu et al., 2023), have emerged to facilitate the implementation of these systems.
In table reasoning, MAS applications remain limited. Existing frameworks like Dater (Ye et al., 2023) employ minimal agent structures tailored to question-answering tasks without extensive collaboration. Similarly, Table-Critic (Yu et al., 2025b) primarily focuses on output refinement rather than interactive reasoning.
Memory Mechanisim for Agent
Memory enables agents to perform coherent, context-aware reasoning across extended tasks. In multi-agent systems, memory serves as the cognitive foundation for retaining observations, decisions, and interaction histories—critical elements for consistent collaboration and adaptation over time (Sumers et al., 2024).
Memory formats determine how content is stored and utilized, with several prominent approaches emerging in recent work. Natural language representations offer semantic richness and interpretability, as demonstrated in Reflexion (Shinn et al., 2023), Voyager (Wang et al., 2023a), and Generative Agents (Park et al., 2023). Vector embeddings enable efficient similarity-based retrieval, a technique central to systems like MemoryBank (Zhong et al., 2023), A-MEM (Xu et al., 2025), and ChatDev (Qian et al., 2024). Meanwhile, structured formats support symbolic reasoning and precise queries, approaches adopted by ChatDB (Hu et al., 2023) and DB-GPT (Zhou et al., 2023b). Beyond storage formats, recent advances have introduced innovative management strategies, including complete interaction storage (Zhong et al., 2023; Modarressi et al., 2024), cache-like designs (Packer et al., 2024), and controller-based architectures (Wang et al., 2025) that dynamically prioritize and maintain relevant information during extended reasoning processes.
Table Reasoning
Research on table reasoning can be broadly classified into fine-tuning-based and prompting-based methods. Fine-tuning approaches like TAPAS (Herzig et al., 2020), Pasta (Gu et al., 2022), TUTA (Wang et al., 2021), and TAPEX (Liu et al., 2022) adapt pre-trained language models to encode table semantics through specialized training objectives. Other works focus on improving alignment between natural language queries and structured data (Eisenschlos et al., 2020; Jiang et al., 2022). Despite their effectiveness, these methods typically require extensive annotated data and feature static reasoning processes without adaptive correction mechanisms.
Prompting-based methods leverage LLMs with minimal training data requirements. Techniques like Chain-of-Thought (Wei et al., 2023), Least-to-Most (Zhou et al., 2023a) and Dater (Ye et al., 2023) perform reasoning by decomposing tasks into explicit steps. Table-specific adaptations include Binder (Cheng et al., 2023), Chain-of-Table (Wang et al., 2024b), ReAcTable (Zhang et al., 2023), and Table-Critic (Yu et al., 2025b), incorporating agent collaboration or ReAct-style reasoning.
Existing table reasoning methods exhibit significant limitations across key design dimensions. While some implement multi-agent architectures or ReAct-based reasoning, none integrates all critical components: dynamic planning, reflection mechanisms, self-refinement, and long-term memory. Our proposed MAPLE  framework addresses these gaps by combining collaborative verification, adaptive planning, and evolving memory structures to achieve more robust and accurate table reasoning.
@@
4:     $m\leftarrow\pi_{\text{archiver-sum}}(\mathcal{M}_{w})$triangleright Distill experience into memory note, Eq. 4
5:     $\mathcal{N}_{\text{archiver}}\leftarrow$ $\pi_{\text{mem-retrieve}}^{\text{archiver}}(m,\mathcal{M}_{l},k,\delta)$triangleright Retrieve $\leq k$ similar memories within $\delta$, Eq. 6
@@
7:         $e\leftarrow\pi_{\text{archiver-evo}}(m,\mathcal{N}_{\text{archiver}})$triangleright Decide evolution actions, Eq. 7
[…]
@@
[…]
@@ B.1 Memory Module
The working memory $(\mathcal{M}_{w})$ temporarily stores all information related to the current task instance, implementing a Shared Message Pool architecture (Hong et al., 2024) for agent communication. As shown in Figure 3, it maintains the original table $\mathcal{T}$, the question $q$, the Solver’s operation history (including intermediate tables and tentative answers), Checker feedback (scores and comments), Reflector analysis (diagnosis and suggestions), and task-level metadata such as the number of remaining attempts.
@@
The long-term memory $(\mathcal{M}_{l})$ captures accumulated knowledge across tasks, supporting continual improvement and experience reuse. Inspired by frameworks like A-MEM (Xu et al., 2025), we adapt their approaches specifically for table-based reasoning challenges.
@@ B.2 Action Module
Following the categorization proposed in CoALA Framework (Sumers et al., 2024), action spaces can be broadly divided into two categories: external actions and internal actions. External actions involve interactions with the external environment, such as controlling a robot, communicating with a human, or navigating a website. As our framework focuses on table reasoning tasks, we do not involve external actions.
@@
Figure 3 summarizes the internal actions associated with each agent in our system. Both the Solver and Checker agents primarily engage in reasoning actions: they read from the working memory, process information according to their designated roles, and write the updated reasoning steps back into the working memory. The Reflector agent performs both reasoning and learning, as it not only updates the working memory but also contributes insights to the long-term memory. The Archiver agent engages in all three action types: retrieval to access relevant experiences, reasoning to analyze current tasks, and learning to evolve the memory base with new knowledge.
@@ B.3 Planning Module
Effective planning is crucial for solving multi-step reasoning tasks, where the sequence and selection of actions can significantly impact final outcomes. Following the categorization proposed by Wang et al. (2024a), planning approaches can be broadly classified into two categories: planning without feedback and planning with feedback.
@@
Environmental Feedback enables the Solver to observe changes in the table state after each operation and decide whether to continue manipulation or derive an answer. Similar to approaches like ReAct (Yao et al., 2023), our framework incorporates thought-action-observation triplets, allowing the Solver to adapt its reasoning trajectory based on real-time observations of how table manipulations affect the environment state. This environmental grounding prevents the accumulation of reasoning errors that plague single-pass methods.
@@ Appendix C Memory Evolution Algorithm
In this section, we present the detailed algorithm for MAPLE’s memory evolution process (Algorithm 2). While the main text describes the conceptual framework and key innovations of our memory system, this appendix provides the complete algorithmic implementation of how new experiences are evaluated, filtered, and integrated into the long-term memory base.
@@ Appendix D Case Study
Solver Round 1 (Section D.1)
@@
Checker Round 1 (Section D.2)
@@
Solver Round 2 (Section D.4D.5)
@@
Checker Round 2 (Section D.6)
@@ D.2.1 User Input
Original Table (Table 5)
@@ D.3.1 User Input
Original Table (Table 5)
@@ D.4.1 User Input
Original Table (Table 5)
@@ D.4.2 LLM Response
  • Intermediate table: see Table 6
@@ D.5.1 User Input
Intermediate Table(Table 6) from Round 4(D.4)
@@ D.6.1 User Input
Original Table (Table 5)
@@ D.7.1 User Input
Original Table (Table 5)

### v2

@@ 4.2 Main Results
Table 2 presents the performance comparison on WikiTQ and TabFact across LLaMA3.3-70B, GPT-4o-mini  and Qwen2.5-72B. Our proposed method, MAPLE, consistently outperforms all baselines across both datasets and model backbones. On WikiTQ, MAPLE  achieves 74.01% and 73.39% accuracy with respective models, while on TabFact, it reaches 90.66% and 86.02%. These results represent substantial gains of up to +8.26% on WikiTQ and +7.41% on TabFact over the strongest baseline methods.
Compared to recent specialized frameworks like Chain-of-Table and ReAcTable, MAPLE  demonstrates consistent improvements across both datasets. The gains are particularly pronounced on WikiTQ (+5.52% with Qwen2.5-72B), aligning with our framework’s strength in handling compositional reasoning tasks that require progressive refinement. For TabFact, the improvements confirm that our approach remains effective even in binary classification settings. For analysis of how table size affects reasoning performance and the impact of multi-round reasoning, see Appendix˜E.
@@ 4.3 Evaluating Agent Contributions
To assess the contribution of each agent in our framework, we conduct an ablation study using the LLaMA3.3-70B model on WiKiTQ. As shown in Table 3, we incrementally introduce agents into the MAPLE  pipeline. Starting from a baseline that directly predicts answers without reasoning, we observe that adding the Solver agent alone leads to a substantial performance boost of +18.2 points, confirming the effectiveness of our ReAct-style multi-step reasoning. Incorporating the Checker further improves accuracy (+2.1), suggesting that verifying answer quality plays a crucial role in reducing erroneous outputs. Introducing the Reflector yields an additional significant gain (+5.2), highlighting the importance of iterative reflection and error correction. Finally, equipping the system with the Archiver enables long-term memory utilization, resulting in peak accuracy of 74.01.
@@ 4.4 Memory Analysis and System Behavior
As $\delta$ increases from 0.3 to 1.0, total memory size decreases dramatically, from 4,078 to 191 notes for WikiTQ and 1,882 to 78 for TabFact. This demonstrates our selective integration approach in action, preventing memory explosion at low thresholds while becoming increasingly selective at higher thresholds. At $\delta=0.3$, nearly all experiences (93.9%) are added to memory, creating potential redundancy, while at $\delta=1.0$, only highly unique experiences (4.4%) are retained, representing two extremes of the memory gradient phenomenon.
@@
Both datasets exhibit remarkably similar memory dynamics despite their different task characteristics, suggesting that these patterns reflect fundamental properties of knowledge organization rather than dataset-specific artifacts.
@@
These findings align with the "approximate learning" theory in cognitive science, which posits that optimal knowledge acquisition occurs when new information is related to existing knowledge that is neither too similar nor too different (10.1093/oxfordhb/9780195376746.013.0042). Our empirical results showing peak evolution at moderate distances ($\delta\approx 0.7$) provide computational evidence for this cognitive principle.
@@ Appendix A Related Work
Model
Multi-Agent
System
External
Tool Use
ReAct Reflection
Self-
Refinement
Long-Term
Memory
Dynamic
Planning
Binder (cheng_binding_2023)
Dater (ye_large_2023)
Chain-of-Table (wang_chain--table_2024)
ReAcTable (zhang_reactable_2023)
Table-Critic (yu2025tablecriticmultiagentframeworkcollaborative)
MAPLE (Ours)
Table 5: Comparison of representative table reasoning methods across key design dimensions.
As illustrated in Table 5, existing table reasoning methods exhibit significant limitations across key design dimensions. While some implement multi-agent architectures or ReAct-based reasoning, none integrates all critical components: dynamic planning, reflection mechanisms, self-refinement, and long-term memory. Our proposed MAPLE  framework addresses these gaps by combining collaborative verification, adaptive planning, and evolving memory structures to achieve more robust and accurate table reasoning.
@@
4:  $m\leftarrow\pi_{\text{archiver-sum}}(\mathcal{M}_{w})$ $\triangleright$ Distill experience into memory note, Eq. 4
5:  $\mathcal{N}_{\text{archiver}}\leftarrow$ $\pi_{\text{mem-retrieve}}^{\text{archiver}}(m,\mathcal{M}_{l},k,\delta)$ $\triangleright$ Retrieve $\leq k$ similar memories within $\delta$, Eq. 6
@@
7:   $e\leftarrow\pi_{\text{archiver-evo}}(m,\mathcal{N}_{\text{archiver}})$ $\triangleright$ Decide evolution actions, Eq. 7
8:   if $e.\texttt{should\_evolve}$ == True then
@@
12:     if $\texttt{update\_neighbor}\in e.\texttt{actions}$ then
@@
15:     $m.\texttt{tags}\leftarrow e.\texttt{tags\_to\_update}$ $\triangleright$ Update semantic tags
16:   end if
@@ B.1 Memory Module
The working memory $(\mathcal{M}_{w})$ temporarily stores all information related to the current task instance, implementing a Shared Message Pool architecture (hong2024metagptmetaprogrammingmultiagent) for agent communication. As shown in Figure 3, it maintains the original table $\mathcal{T}$, the question $q$, the Solver’s operation history (including intermediate tables and tentative answers), Checker feedback (scores and comments), Reflector analysis (diagnosis and suggestions), and task-level metadata such as the number of remaining attempts.
@@
The long-term memory $(\mathcal{M}_{l})$ captures accumulated knowledge across tasks, supporting continual improvement and experience reuse. Inspired by frameworks like A-MEM (xu2025amemagenticmemoryllm), we adapt their approaches specifically for table-based reasoning challenges.
@@ B.2 Action Module
Following the categorization proposed in CoALA Framework (sumers2024cognitivearchitectureslanguageagents), action spaces can be broadly divided into two categories: external actions and internal actions. External actions involve interactions with the external environment, such as controlling a robot, communicating with a human, or navigating a website. As our framework focuses on table reasoning tasks, we do not involve external actions.
@@
Figure˜3 summarizes the internal actions associated with each agent in our system. Both the Solver and Checker agents primarily engage in reasoning actions: they read from the working memory, process information according to their designated roles, and write the updated reasoning steps back into the working memory. The Reflector agent performs both reasoning and learning, as it not only updates the working memory but also contributes insights to the long-term memory. The Archiver agent engages in all three action types: retrieval to access relevant experiences, reasoning to analyze current tasks, and learning to evolve the memory base with new knowledge.
@@ B.3 Planning Module
Effective planning is crucial for solving multi-step reasoning tasks, where the sequence and selection of actions can significantly impact final outcomes. Following the categorization proposed by Wang_2024, planning approaches can be broadly classified into two categories: planning without feedback and planning with feedback.
@@
Environmental Feedback enables the Solver to observe changes in the table state after each operation and decide whether to continue manipulation or derive an answer. Similar to approaches like ReAct (yao2023reactsynergizingreasoningacting), our framework incorporates thought-action-observation triplets, allowing the Solver to adapt its reasoning trajectory based on real-time observations of how table manipulations affect the environment state. This environmental grounding prevents the accumulation of reasoning errors that plague single-pass methods.
@@ Appendix C Memory Evolution Algorithm
In this section, we present the detailed algorithm for MAPLE’s memory evolution process (Algorithm˜2). While the main text describes the conceptual framework and key innovations of our memory system, this appendix provides the complete algorithmic implementation of how new experiences are evaluated, filtered, and integrated into the long-term memory base. Specifically, the AddLinks function corresponds to the Strengthen option in Figure˜11: it does not alter the metadata of the neighboring memories but instead appends their memory_id to the Links field of the current memory. In contrast, the UpdateMetadata function corresponds to the Update_Neighbor option in Figure˜11: it summarizes new tags and context based on both the current and neighboring memories, and updates these fields in each memory accordingly.
@@ Appendix D Case Study
Solver Round 1 (Section˜D.1)
@@
Checker Round 1 (Section˜D.2)
@@
Solver Round 2 (Section˜D.4D.5)
@@
Checker Round 2 (Section˜D.6)
@@ D.2.1 User Input
Original Table (Table˜6)
@@ D.3.1 User Input
Original Table (Table˜6)
@@ D.4.1 User Input
Original Table (Table˜6)
@@ D.4.2 LLM Response
  • Intermediate table: see Table˜7
@@ D.5.1 User Input
Intermediate Table(Table˜7) from Round 4(D.4)
@@ D.6.1 User Input
Original Table (Table˜6)
@@ D.7.1 User Input
Original Table (Table˜6)
@@ E.2 Impact of Multi-Round Reasoning
E.3 Impact of Memory Evolution Strategies
Dataset
Evolve Type
Memory
Count
Memory
Ratio (%)
Evolution
Count
Evolution
Ratio (%)
# Evolved
Memories
Evolution
Efficiency ↑
Med. Strengthen
Distance
Med. Update
Distance
Accuracy
WiKiTQ Always Evolve 1023 23.5% 711 69.5% 936 1.32 0.64 0.64 71.42
Evolve Every $n$ Entries 1023 23.5% 147 14.4% 190 1.29 0.64 0.64 69.96
LLM-Based 1023 23.5% 667 65.2% 820 1.23 0.64 0.63 74.01
Never Evolve 1023 23.5% 0 00.0% 0 0.00 NA NA 67.89
TabFact Always Evolve 427 21.1% 337 78.9% 423 1.26 0.61 0.61 85.84
Evolve Every $n$ Entries 427 21.1% 68 15.9% 90 1.32 0.63 0.63 85.75
LLM-Based 427 21.1% 319 74.7% 372 1.17 0.60 0.58 90.66
Never Evolve 427 21.1% 0 00.0% 0 0.00 NA NA 86.29
Table 8: Performance of different memory evolution strategies under fixed parameters (similarity threshold $\delta$=0.7, update interval n=5) on the WiKiTQ  and TabFact datasets.
Table˜8 compares different memory evolution strategies under fixed parameters ($\delta$=0.7, n=5) on the WiKiTQ  and TabFact datasets. The results show that memory evolution has a substantial impact on QA performance. In both datasets, the Never Evolve setting yields the lowest accuracy (67.89% on WiKiTQ  and 86.29% on TabFact), confirming that static memory limits the system’s adaptability. In contrast, Always Evolve improves performance, but its brute-force nature leads to a large number of updates (69.5% and 78.9% evolution ratios on WiKiTQ  and TabFact, respectively), introducing redundancy without fully translating into accuracy gains. Similarly, the periodic policy (Evolve Every n Entries) is overly rigid, resulting in moderate efficiency but still lower accuracy than more adaptive approaches.
The LLM-Based strategy consistently achieves the best accuracy (74.01% on WiKiTQ  and 90.66% on TabFact) while maintaining a balanced evolution efficiency. Unlike naive policies, it selectively evolves memory entries guided by semantic signals, allowing the system to capture meaningful updates while avoiding unnecessary changes. This flexibility explains why its accuracy surpasses both the always-update and fixed-frequency baselines, despite fewer total evolutions. These findings highlight that LLM-guided memory evolution is effective, striking a balance between adaptability and efficiency, and avoiding the rigidity of rule-based strategies.